The population of Johnstown is increasing linearly in time. In 2000, the population was 47. In 2006, it was 77. Let y be the populatiion and x be the year using 0 for the year 2000.
Write an equation representing this relationship.
Estimate the population in 2018.
Any advise is great appreciated!

Answer by mananth(16949)

The population of Johnstown is increasing linearly in time. In 2000, the population was 47. In 2006, it was 77. Let y be the populatiion and x be the year using 0 for the year 2000.
Write an equation representing this relationship.
Estimate the population in 2018.
...
y2000 47
y2006 77
...
increase per year = 30/6 = 5
...
y= 47+5x
...
y 2018
y=47+5*18
y=235+18
y=253
...
m.ananth@hotmail.ca